# Minimal pole representation for spectral functions

### v1

@@ I Introduction
The accurate representation of a continuous quantum system by a few discrete degrees of freedom is a fundamental problem in condensed matter science. Applications of this problem appear whenever the solution of quantum impurity problems requires a Hamiltonian representation [1, 2, 3, 4] with a few degrees of freedom [1, 5, 6, 7, 8]; whenever transport in open systems is modeled by coupling dissipative bath states to a correlated subsystem [9, 10, 11], and wherever fast quasi-analytic calculations in a Green’s function language are needed [12].
Compact and accurate representations of Green’s functions are essential in computational many-body simulations [13, 14, 15, 16, 17, 18, 19, 20]. A compact and accurate representation respecting the analytic properties of equilibrium Green’s function has recently been introduced [21, 22] in the context of analytic continuation [23, 24, 25, 26, 27, 28, 29] of imaginary-frequency Matsubara Green’s functions and is based on the realization that most objects of interest in many-body theory, such as retarded or advanced Green’s functions, susceptibilities, and spectral functions are related to Nevanlinna functions [25, 26]. The retarded Green’s function in particular is analytic in the upper half of the complex plane, its imaginary part coincides (up to a constant prefactor) with the spectral function just above the real axis, and all of its poles lie in the lower half of the complex plane. With the help of suitable holomorphic mappings, the problem of accurately approximating the Green’s function in the upper half plane simplifies to an exponential approximation problem [21, 22, 28, 27], for which well-established signal processing methodology exists [30, 31, 32].
Although functions known directly on the real axis do not suffer from the ill-conditioning of the analytic continuation problem, in practical applications they, too, benefit from systematically improvable compact representations in terms of poles in the complex plane. Such representations enable fast calculations of diagrams in perturbation theory [12, 33], fast propagation of real-time quantities in hierarchical equation of motion calculations [9, 10, 11], and a systematic analysis of the features contained in spectral data [21, 22].
@@ II Method
with $z\in\mathbb{C}$ and ${A}(z=\omega)\approx A^{\text{ex}}(\omega)$, where $A_{l}^{(\mathrm{dn})}$ and $\xi_{l}^{(\mathrm{dn})}$ are complex weights and pole locations in the lower half of the complex plane, while $A_{l}^{(\mathrm{up})}$ and $\xi_{l}^{(\mathrm{up})}$ are their counterparts in the upper half plane.
@@ II.1 Minimal Pole Method
As first discussed in the context of analytic continuation in Refs. [27, 28], applying a holomorphic mapping to Matsubara data followed by the application of the residue theorem reformulates the continuation problem into an exponential approximation or ‘Prony’ problem. More recently, Refs. [21, 22] demonstrated that once a continuous imaginary interval is adequately approximated from discrete Matsubara data, this methodology ensures systematic convergence of the spectral function with increasing data precision. The spectral approximation problem on the real axis can be formulated in much the same way, with the main difference between the analytic continuation and the real-frequency problem being the holomorphic mappings. In this paper, we abbreviate the minimal pole method as MPM.
Among the infinitely many possible choices, we find the following two holomorphic mappings particularly useful for the real-frequency problem: When $A(\omega)$ has spectral weight at all frequencies, we propose to apply the Möbius transform
@@
In the case where $A(\omega)$ has compact support in the interval $[\omega_{\rm min},\omega_{\rm max}]$, we propose to apply the transform
@@
where $\omega_{\rm m}=(\omega_{\rm min}+\omega_{\rm max})/2$, $\Delta\omega_{\rm h}=(\omega_{\rm max}-\omega_{\rm min})/2$, and the branch of the square root is chosen such that $|u|\leq 1$. We denote this mapping as M2. Illustrations for both M1 and M2 are shown in Fig. 1.
@@
where $A(z)\equiv A^{\prime}(u)$. Since the function is known on the unit circle of the mapped plane, one may apply the residue theorem to extract the location of the mapped poles via contour integrals along the unit circle. As shown in Refs. [27, 28], after defining the contour integral $h_{k}$ as
@@
Given real-axis spectral functions, the integrals $h_{k}$ ($k=0,1,2...$) can be numerically calculated to high precision using standard quadrature rules. The computation is terminated when the target cutoff precision $\varepsilon$ is reached, i.e., when $|h_{k}|\leq\varepsilon$, or when a maximum number of terms $k_{\rm max}$ are computed, whichever occurs first.
Knowing $h_{k}$, the Prony-type approximation problem of Eq. (11) can then be solved with standard solution methods for the Prony problem, resulting in $M$ values $A_{l}^{\prime\rm(dn)}$ and $x_{l}^{\prime\rm(dn)}$. $M$ is either determined a priori as a fixed number of poles, or adjusted dynamically such that a target precision of less than $\varepsilon$ is reached. Appendix A contains more details for the Prony problem and its solution with the ESPRIT method.
Finally, the values $x_{l}^{\rm(dn)}$ are recovered using the inverse transform [27, 28]
@@
and $A_{l}^{\rm(dn)}$ are obtained using the relation [21, 22]
@@
In the case of multi-orbital systems and matrix-valued retarded Green’s functions, it is often physically insightful to approximate the system using shared complex poles with matrix-valued weights. In this case, a matrix-valued generalization of the ESPRIT algorithm [22] can be employed.
@@ II.2.1 From the spectral function to the Green’s function on the complex plane
The Green’s function $G(z)$ in the frequency domain is related to the spectral function $A(\omega)$ by [34, 35, 36, 37, 38]
@@
$G(\omega+i0^{+})=-2\pi i\sum_{l}\frac{A_{l}^{({\rm dn})}}{\omega+i0^{+}-\xi_{l}^{({\rm dn})}}.$ (26)
@@
As we will show, $A(\omega)$ can be approximated by complex poles to arbitrary precision. By Eq. (26), the approximation of $G(z)$ at $z=\omega+i0^{+}$ ($\omega\in(-\infty,+\infty)$) can also be achieved to arbitrary precision. Since both the exact Green’s function in Eq. (20) and the approximation in Eq. (27) are analytic for ${\rm Im}\,z>0$ and their differences on the boundary can be made arbitrarily small, the Green’s function throughout the upper half-plane can be fully described by ${A_{l}^{({\rm dn})}}$ and $\xi_{l}^{({\rm dn})}$ with arbitrary precision according to the maximum modulus principle. This confirms that the complex pole representation is numerically sufficient to capture all features of $G(z)$, which forms the theoretical foundation for works based on the complex pole representation, such as Refs. [27, 28, 21, 22, 42, 43].
@@ II.2.2 Decomposition of exponentials for real-time evolution
In the simulation of open quantum systems, the bath correlation function (BCF), denoted as $C(t)$, plays a central role as it fully characterizes the influence of the environment on the system of interest. Simulation methods, such as the hierarchical equations of motion (HEOM) [9, 11, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53], the hierarchy of pure states [54], the pseudomodes approach [55], auxiliary master equation methods (AMEA) [10, 56, 57], and the recently developed Quasi-Lindblad theory [58], rely on its efficient expansion of exponentials,
@@ III.1 Approximation of typical spectral functions
We evaluate the pole approximation method by examining three paradigmatic spectral functions that are analytically known on the real axis: a Kondo-like structure, a bosonic multi-orbital spectral function, and a semicircular density of states for MPM (M1). The free parameter $\omega_{p}$ in the mapping is set to $\omega_{p}=2$. For the calculation of the contour integral, we use $k_{\rm max}=3000$ and a precision of $\varepsilon=10^{-12}$, as defined in Sec. II.1. MPM (M2) yields very similar results, except that the approximation is confined to the finite interval $[\omega_{\rm min},\omega_{\rm max}]$ (not shown).
@@
which corresponds to the local density of states for an infinite-dimensional Bethe lattice [59]. The hopping parameter is set to $t=1$, resulting in square-root singularities at $\omega=\pm 2$. As shown in Fig. 2(k), the approximation converges exponentially away from the singularities as $M$ increases. However, as depicted in Fig. 2(j), the presence of singularities slows down convergence. Despite the challenges posed by these singularities, their impact on practical calculations based on complex poles is typically small, as will be demonstrated in the next section. Fig. 2(l) shows the distribution of poles which cluster around the singular points.
@@ III.3 Decomposing model spectral densities for real-time evolution
In this section, we evaluate the performance of the methods on the BCF derived from a power-law spectral density with an exponential cutoff. We compare our results with those obtained using the adaptive Antoulas–Anderson (AAA) algorithm [60, 61], which operates on the real-frequency domain, as well as with those obtained from the direct decomposition of the BCF using ESPRIT [32, 62].
@@
and extended to negative frequencies $\omega<0$ by $J(\omega)=-J(-\omega)$. This form is classified as Ohmic, sub-Ohmic, or super-Ohmic for $s=1$, $10<s<$ 1, and $s>1$, respectively. The corresponding BCF has the analytic form [63, 64]
@@
In this section, we fix the parameters to $\alpha=1$, $\omega_{c}=50\;{\rm cm}^{-1}$, and $t_{c}=2000\;{\rm fs}$, following Ref. [53]. The sampling points for ESPRIT are chosen on a uniform grid,
@@
whereas for AAA, a logarithmic discretization [65, 53] over $[-\omega_{\rm max},\omega_{\rm max}]$ is employed to ensure efficient sampling in the low-frequency domain:
@@
In our simulations, both $N_{t}$ and $N_{\omega}$ are fixed at 5000. We use $\omega_{p}=0.1\;{\rm fs}^{-1}$ for MPM (M1) and $\omega_{\rm max}=-\omega_{\rm min}=0.5\;{\rm fs}^{-1}$ for MPM (M2) consistently. Additionally, $k_{\rm max}$ is fixed at 1000 for both MPM (M1) and MPM (M2) throughout this section.
@@ III.3.1 Ohmic bath
Figure 4: Comparison of AAA (red), ESPRIT (blue), MPM (M1) (green), and MPM (M2) (orange) for an Ohmic bath at $T=300\,{\rm K}$. Left panel: (a) Exact effective spectral function $A(\omega)$ and (d) exact bath correlation function $C(\omega)$. Middle panel: Relative error of (b) $A(\omega)$ as defined in Eq. (41) and (e) $C(t)$ as defined in Eq. (42). All approximations use $M=8$ terms. The vertical line separates the inside and outside of the approximated real-time interval. The relative difference $|\hat{C}^{(L_{1})}(t)-\hat{C}^{(L_{2})}(t)|/|\hat{C}^{(L_{1})}(0)|$ is shown by the thick gray curve. Right panel: Mean error (Eqs. (43) and (44)) as a function of $M$, evaluated (c) within the interval $[0,t_{c}]$ and (f) over the long tail $[t_{c},10t_{c}]$.
We begin our discussion with the Ohmic bath ($s=1$) at a temperature of $T=300\;{\rm K}$. As a reference, the left column of Fig. 4 displays the exact input function. In the middle panel, we fix $M=8$ for all methods and compare their performance in both the frequency and time domains.
@@
We now examine the same system at a much lower temperature, $T=0.001\;{\rm K}$. In this regime, the effective spectrum varies rapidly near the origin. As shown in Fig. 5, a convergence slowdown as a function of $M$ is observed at low temperature for the methods discussed in this paper. Their performance trends remain consistent with those observed at higher temperatures. Although AAA occasionally achieves an accuracy in the long tail similar to MPM (M1), its convergence behavior is less robust and its accuracy within the finite interval is consistently lower than that of MPM (M1).
Figure 5: Same setup as Fig. 4, except that temperature is now set to $T=0.001\,{\rm K}$. Top panel: (a) Exact spectral function. (b) Exact bath correlation function. Bottom panel: Mean error for (c) the approximated interval and (d) the long-tail.

### v2

@@ I Introduction
The accurate representation of a continuous quantum system by a few discrete degrees of freedom is a fundamental problem in condensed matter science. Applications of this problem appear whenever the solution of quantum impurity problems requires a Hamiltonian representation Caffarel and Krauth (1994); Zgid, Gull, and Chan (2012); Shee and Zgid (2019); Zhu et al. (2019) with a few degrees of freedom Caffarel and Krauth (1994); Koch, Sangiovanni, and Gunnarsson (2008); Sénéchal (2010); Liebsch and Ishida (2011); Mejuto-Zaera et al. (2020); whenever transport in open systems is modeled by coupling dissipative bath states to a correlated subsystem Tanimura and Kubo (1989); Arrigoni, Knap, and von der Linden (2013); Tanimura (2020), and wherever fast quasi-analytic calculations in a Green’s function language are needed Gazizova et al. (2024).
Compact and accurate representations of Green’s functions are essential in computational many-body simulations Boehnke et al. (2011); Kananenka et al. (2016); Shinaoka et al. (2017, 2018); Li et al. (2020); Kaye and Golež (2021); Dong, Gull, and Strand (2022); Kaye, Chen, and Parcollet (2022). A compact and accurate representation respecting the analytic properties of equilibrium Green’s function has recently been introduced Zhang and Gull (2024); Zhang, Yu, and Gull (2024) in the context of analytic continuation Jarrell and Gubernatis (1996); Burnier and Rothkopf (2013); Fei, Yeh, and Gull (2021); Fei et al. (2021); Ying (2022a, b); Huang, Gull, and Lin (2023) of imaginary-frequency Matsubara Green’s functions and is based on the realization that most objects of interest in many-body theory, such as retarded or advanced Green’s functions, susceptibilities, and spectral functions are related to Nevanlinna functions Fei, Yeh, and Gull (2021); Fei et al. (2021). The retarded Green’s function in particular is analytic in the upper half of the complex plane, its imaginary part coincides (up to a constant prefactor) with the spectral function just above the real axis, and all of its poles lie in the lower half of the complex plane. With the help of suitable holomorphic mappings, the problem of accurately approximating the Green’s function in the upper half plane simplifies to an exponential approximation problem Zhang and Gull (2024); Zhang, Yu, and Gull (2024); Ying (2022b, a), for which well-established signal processing methodology exists de Prony (1795); Hua and Sarkar (1990); Roy and Kailath (1989).
Although functions known directly on the real axis do not suffer from the ill-conditioning of the analytic continuation problem, in practical applications they, too, benefit from systematically improvable compact representations in terms of poles in the complex plane. Such representations enable fast calculations of diagrams in perturbation theory Gazizova et al. (2024); Kaye et al. (2024), fast propagation of real-time quantities in hierarchical equation of motion calculations Tanimura and Kubo (1989); Arrigoni, Knap, and von der Linden (2013); Tanimura (2020), and a systematic analysis of the features contained in spectral data Zhang and Gull (2024); Zhang, Yu, and Gull (2024).
@@ II Method
with $z\in\mathbb{C}$ and ${A}(z=\omega)\approx A^{\text{ex}}(\omega)$, where $A_{l}^{(\mathrm{dn})}$ and $\xi_{l}^{(\mathrm{dn})}$ are complex weights and pole locations in the lower half of the complex plane, while $A_{l}^{(\mathrm{up})}$ and $\xi_{l}^{(\mathrm{up})}$ are their counterparts in the upper half-plane.
@@ II.1 Minimal Pole Method
As first discussed in the context of analytic continuation in Refs. Ying (2022a, b), applying a holomorphic mapping to Matsubara data followed by the application of the residue theorem reformulates the continuation problem into an exponential approximation or ‘Prony’ problem. More recently, Refs. Zhang and Gull (2024); Zhang, Yu, and Gull (2024) demonstrated that once a continuous imaginary interval is adequately approximated from discrete Matsubara data, this methodology ensures systematic convergence of the spectral function with increasing data precision. The spectral approximation problem on the real axis can be formulated in much the same way, with the main difference between the analytic continuation and the real-frequency problem being the holomorphic mappings. In this paper, we abbreviate the minimal pole method as MPM.
Among the infinitely many possible choices, we find the following two holomorphic mappings particularly useful for the real-frequency problem: one is used to construct a controlled approximation over the entire real-frequency axis $\omega\in(-\infty,+\infty)$, while the other is designed for controlled approximation over a finite interval $\omega\in[\omega_{\min},\omega_{\max}]$ of interest. We note that other mappings are possible.
For the former case, which takes the spectral weight of $A(\omega)$ over the entire frequency range as input, we propose applying the Möbius transform
@@
For the latter case, where $A(\omega)$ has compact support in the interval $[\omega_{\rm min},\omega_{\rm max}]$, we propose applying the transform
@@
where $\omega_{\rm m}=(\omega_{\rm min}+\omega_{\rm max})/2$, $\Delta\omega_{\rm h}=(\omega_{\rm max}-\omega_{\rm min})/2$, and the branch of the square root is chosen such that $|u|\leq 1$. We denote this mapping as M2. Illustrations of both M1 and M2 are shown in Fig. 1, with a detailed performance comparison provided in Sec. III.
@@
where $A(z)\equiv A^{\prime}(u)$. Since the function is known on the unit circle of the mapped plane, one may apply the residue theorem to extract the location of the mapped poles via contour integrals along the unit circle. As shown in Refs. Ying (2022a, b), after defining the contour integral $h_{k}$ as
@@
Given spectral functions on the real axis, the integrals $h_{k}$ ($k=0,1,2,\ldots$) can be numerically evaluated to high precision using standard quadrature rules. Since the poles $\{A_{l}^{\prime},\xi_{l}^{\prime}\}$ lie strictly inside the unit circle, $h_{k}$ decays with increasing $k$ according to Eq. (11). Therefore, the computation of $h_{k}$ can be terminated once the target precision $\varepsilon$ is reached—i.e., when $|h_{k}|\leq\varepsilon$—or when a maximum number of terms $k_{\rm max}$ is reached if $h_{k}$ decays slowly, whichever occurs first.
Knowing $h_{k}$, the Prony-type approximation problem of Eq. (11) can then be solved with standard solution methods for the Prony problem, resulting in $M$ values $A_{l}^{\prime\rm(dn)}$ and $x_{l}^{\prime\rm(dn)}$. $M$ is either determined a priori as a fixed number of poles, or adjusted dynamically such that a target precision is reached. Appendix A contains more details for the Prony problem and its solution with the ESPRIT method.
Finally, the values $x_{l}^{\rm(dn)}$ are recovered using the inverse transform Ying (2022a, b)
@@
and $A_{l}^{\rm(dn)}$ are obtained using the relation Zhang and Gull (2024); Zhang, Yu, and Gull (2024)
@@
In the case of multi-orbital systems and matrix-valued retarded Green’s functions, it is often physically insightful to approximate the system using shared complex poles with matrix-valued weights. In this case, a matrix-valued generalization of the ESPRIT algorithm Zhang, Yu, and Gull (2024) can be employed.
@@ II.2.1 From the spectral function to the Green’s function on the complex plane
The Green’s function $G(z)$ in the frequency domain is related to the spectral function $A(\omega)$ by Smit (2022); Negele (2018); Mahan (2013); Stefanucci and van Leeuwen (2013); Coleman (2015)
@@
$G(\omega+i0^{+})=-2\pi i\sum_{l=1}^{M}\frac{A_{l}^{({\rm dn})}}{\omega+i0^{+}-\xi_{l}^{({\rm dn})}}.$ (26)
@@
As we will show, $A(\omega)$ can be approximated by complex poles to arbitrary precision over $\omega\in(-\infty,+\infty)$ by utilizing the M1 mapping. By Eq. (26), the approximation of $G(z)$ at $z=\omega+i0^{+}$ ($\omega\in(-\infty,+\infty)$) can also be achieved to arbitrary precision. Since both the exact Green’s function in Eq. (20) and the approximation in Eq. (27) are analytic for ${\rm Im}\,z>0$ and their differences on the boundary can be made arbitrarily small, the Green’s function throughout the upper half-plane can be fully described by ${A_{l}^{({\rm dn})}}$ and $\xi_{l}^{({\rm dn})}$ with arbitrary precision according to the maximum modulus principle. This confirms that the complex pole representation is numerically sufficient to capture all features of $G(z)$, which forms the theoretical foundation for works based on the complex pole representation, such as Refs. Ying (2022a, b); Zhang and Gull (2024); Zhang, Yu, and Gull (2024); Goswami, Barros, and Carbone (2024); Huang and Yue (2025).
@@ II.2.2 Decomposition of exponentials for real-time evolution
In the simulation of open quantum systems, the bath correlation function (BCF), denoted as $C(t)$, plays a central role as it fully characterizes the influence of the environment on the system of interest. Simulation methods, such as the hierarchical equations of motion (HEOM) Tanimura and Kubo (1989); Tanimura (2020, 1990); Ishizaki and Tanimura (2005); Hu, Xu, and Yan (2010); Hu et al. (2011); Duan et al. (2017); Erpenbeck et al. (2018); Cui et al. (2019); Xu et al. (2022); Chen et al. (2022); Takahashi et al. (2024), the hierarchy of pure states Suess, Eisfeld, and Strunz (2014), the pseudomodes approach Tamascelli et al. (2018), auxiliary master equation methods (AMEA) Arrigoni, Knap, and von der Linden (2013); Dorda et al. (2014); Werner, Lotze, and Arrigoni (2023), and the recently developed Quasi-Lindblad theory Park et al. (2024), rely on its efficient expansion of exponentials,
@@
A similar procedure can be readily applied in the zero-temperature limit. In this case, Eq. (II.2.2) simplifies to
[…]
One can then define the effective spectral function as $A(\omega)=2J(\omega)$ for $\omega>0$, and $A(\omega)=0$ otherwise. The relation in Eq. (30) remains valid in this limit.
@@ III.1 Approximation of typical spectral functions
We evaluate the pole approximation method by examining three paradigmatic spectral functions that are analytically known on the real axis: a Kondo-like structure, a bosonic multi-orbital spectral function, and a semicircular density of states for MPM (M1). For convenience, we adopt atomic units in Secs. III.1 and III.2, as is customary in studies of model Hamiltonians within the condensed matter community. The free parameter $\omega_{p}$ in the mapping is set to $\omega_{p}=2$. For the calculation of the contour integral, we use $k_{\rm max}=3000$ and a precision of $\varepsilon=10^{-12}$, as defined in Sec. II.1. MPM (M2) yields very similar results, except that the approximation is confined to the finite interval $[\omega_{\rm min},\omega_{\rm max}]$ (not shown).
@@
which corresponds to the local density of states for an infinite-dimensional Bethe lattice Metzner and Vollhardt (1989). The hopping parameter is set to $t=1$, resulting in square-root singularities at $\omega=\pm 2$. As shown in Fig. 2(k), the approximation converges exponentially away from the singularities as $M$ increases. However, as depicted in Fig. 2(j), the presence of singularities slows down convergence. Despite the challenges posed by these singularities, their impact on practical calculations based on complex poles is typically small, as will be demonstrated in the next section. Fig. 2(l) shows the distribution of poles which cluster around the singular points.
@@ III.3 Decomposing model spectral densities for real-time evolution
In this section, we evaluate the performance of the methods on the BCF derived from a power-law spectral density with an exponential cutoff. To better serve the open quantum systems community, we explicitly provide the units of physical quantities in Secs. III.3 and III.4. We compare our results with those obtained using the adaptive Antoulas–Anderson (AAA) algorithm Nakatsukasa, Sète, and Trefethen (2018); Nakatsukasa and Trefethen (2020), which operates on the real-frequency domain, as well as with those obtained from the direct decomposition of the BCF using ESPRIT Roy and Kailath (1989); Potts and Tasche (2013).
@@
and extended to negative frequencies $\omega<0$ by $J(\omega)=-J(-\omega)$. This form is classified as Ohmic, sub-Ohmic, or super-Ohmic for $s=1$, $10<s<$ 1, and $s>1$, respectively. The corresponding BCF has the analytic form Brandes (2003); Weiss (2012)
@@
In this section, we fix the parameters to $\alpha=1$, $\omega_{c}=50\;{\rm cm}^{-1}\approx 9.4\times 10^{-3}\;{\rm fs}^{-1}$, and $t_{c}=2000\;{\rm fs}$, following Ref. Takahashi et al. (2024). The sampling points for ESPRIT are chosen on a uniform grid,
@@
whereas for AAA, a logarithmic discretization Walters, Allen, and Makri (2017); Takahashi et al. (2024) over $[-\omega_{\rm max},\omega_{\rm max}]$ is employed to ensure efficient sampling in the low-frequency domain:
@@
In our simulations, both $N_{t}$ and $N_{\omega}$ are fixed at 5000. We consistently use $\omega_{p}=0.1\;{\rm fs}^{-1}\approx 10.6\;\omega_{c}$ for MPM (M1), and $\omega_{\rm max}=-\omega_{\rm min}=0.5\;{\rm fs}^{-1}=1000\;t_{c}^{-1}\approx 53.1\;\omega_{c}$ for MPM (M2), which matches the value used in AAA. Additionally, $k_{\rm max}$ is fixed at 1000 for both MPM (M1) and MPM (M2) throughout this section. We note that the choice of $\omega_{p}$ in MPM (M1) is independent of $t_{c}$; instead, it is determined by the shape of $A(\omega)$. In contrast, we observe that for such systems, $\omega_{\rm min}$ and $\omega_{\rm max}$ in MPM (M2) are inversely proportional to the time cutoff.
@@ III.3.1 Ohmic bath
Figure 4: Comparison of AAA (red), ESPRIT (blue), MPM (M1) (green), and MPM (M2) (orange) for an Ohmic bath at $T=300\,{\rm K}$ ($\beta\omega_{c}\approx 0.240$). Left panel: (a) Exact effective spectral function $A(\omega)$ and (d) exact bath correlation function $C(\omega)$. Middle panel: Relative error of (b) $A(\omega)$ as defined in Eq. (42) and (e) $C(t)$ as defined in Eq. (43). All approximations use $M=8$ terms. The vertical line separates the inside and outside of the approximated real-time interval. The relative difference $|\hat{C}^{(L_{1})}(t)-\hat{C}^{(L_{2})}(t)|/|\hat{C}^{(L_{1})}(0)|$ is shown by the thick gray curve. Right panel: Mean error (Eqs. (44) and (45)) as a function of $M$, evaluated (c) within the interval $[0,t_{c}]$ and (f) over the long tail $[t_{c},10t_{c}]$.
We begin our discussion with the Ohmic bath ($s=1$) at a temperature of $T=300\;{\rm K}$, corresponding to an inverse temperature of $\beta\approx 25.5\;{\rm fs}$. As a reference, the left column of Fig. 4 displays the exact input function. In the middle panel, we fix $M=8$ for all methods and compare their performance in both the frequency and time domains.
@@
We now examine the same system at a much lower temperature, $T=0.001\;{\rm K}$, corresponding to an inverse temperature of $\beta\approx 7.64\times 10^{6}\;{\rm fs}$. In this regime, the effective spectrum varies rapidly near the origin. As shown in Fig. 5, no convergence slowdown with respect to $M$ is observed at low temperature for the methods discussed in this paper. Their performance trends remain consistent with those observed at higher temperatures. Although AAA occasionally achieves an accuracy in the long tail similar to MPM (M1), its convergence behavior is less robust and its accuracy within the finite interval is consistently lower than that of MPM (M1).
Figure 5: Same setup as Fig. 4, except that temperature is now set to $T=0.001\,{\rm K}$ ($\beta\omega_{c}\approx 7.19\times 10^{4}$). Top panel: (a) Exact spectral function. (b) Exact bath correlation function. Bottom panel: Mean error for (c) the approximated interval and (d) the long-tail.